Adv. Math. Econ. 5, 153-166 (2003)

Monte Carlo method for pricing of Bermuda type derivatives

Shigeo Kusuoka

Graduate School of Mathematical Sciences, The University of Tokyo, 3-8-1 Komaba, Meguro-ku, Tokyo 153-8914, Japan

Received: August 2, 2002 JEL classification: C63, G12 Mathematical Subject Classification (2000): 65C05, 60G40 Abstract. The author gives a rigorous justification of a certain Monte Carlo Method for pricing of Bermuda type derivatives given by Longstaff-Schwartz. Key words: Mathematical finance, Monte Carlo method